OpenStudy (anonymous):

express in simplest form 4x^2/x^2+7x+12/2x-4/x+3 asap!!

OpenStudy (anonymous):

The solution to this question would be: \[7x + 7 = \frac{ 2 }{ x }\]

OpenStudy (anonymous):

can you tell me how you got that?

OpenStudy (moss65):

Actually, it is 7+7x+2/x

OpenStudy (anonymous):

okay but how did you get that answer??

OpenStudy (anonymous):

If your expressing this equation in simplest form then it's:\[7x + 7 = \frac{ 2 }{ x }\]

OpenStudy (anonymous):

I did this stuff all through Middle School and currently, I'm taking college courses. I'm experienced in this type of stuff.

OpenStudy (anonymous):

i just dont know how youre coming up with this answer

OpenStudy (moss65):

4(x^2/x^2)=4 12/2x=6/x 6/x-4/x=2/x 4+7x+2/x+3=7+7x+2/x